40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS IMG_3942.JPG
    Reply  Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS IMG_7471.JPGIMG_7472.JPGIMG_7473.JPGIMG_7474.JPG
       Reply  Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS IMG_7476.JPGIMG_7477.JPGIMG_7478.JPGIMG_7479.JPG
          Reply  Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled 
             Reply  Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement IMG_7497.JPGFOL_schematic.pdf20170825_IR_ALS.pdf
                Reply  Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened IMG_7605.JPG
                Reply  Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement 
                      Reply  Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement IR_ALS_noise.pdf
Message ID: 13254     Entry time: Fri Aug 25 15:54:14 2017     In reply to: 13246     Reply to this: 13255   13257
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS noise measurement 

[Kira, gautam]

Attachment #1 - Photo of the revamped beat setup. The top panel has to be installed. New features include:

  • Regulated power supply via D1000217.
  • Single power switch for both PDs.
  • Power indicator LED.
  • Chassis ground isolated from all other electronic grounds. For this purpose, I installed all the elctronics on a metal plate which is only connected to the chassis via nylon screws. The TO220 package power regulator ICs have been mounted with the TO220 mounting kits that provide a thin piece of plastic that electrically insulates its ground from the chassis ground.
  • PD outputs routed through 20dB coupler on front panel for diagnostic purposes.
  • Fiber routing has been cleaned up a little. I installed a winding fixture I got from Johannes, but perhaps we can install another one of these on top of the existing one to neaten up the fiber layout further.
  • 90-10 light splitter (meant for diagnostic purposes) has been removed because of space constraints. 

Attachment #2 - Power budget inside the box. Some of these FC/APC connectors seem to not offer good coupling between the two fibers. Specifically, the one on the front panel meant to accept the PSL light input fiber seems particularly bad. Right now, the PSL light is entering the box through one of the front panel connectors marked "PSL + X out". I've also indicated the beat amplitude measured with an RF analyzer. Need to do the math now to confirm if these match the expected amplitudes based on the power levels measured.

Attachment #3 - We repeated the measurement detailed here. The X arm (locked to IR) was used for this test. The "X" delay line electronics were connected to the X green beat PD, while the "Y" delay line electronics were connected to the X IR beat PD. I divided the phase tracker Hz calibration factor by 2 to get IR Hz for the Y arm channels. IR beat was at ~38MHz, green beat was at ~76MHz. The broadband excess noise seen in the previous test is no longer present. Indeed, below ~20Hz, the IR beat seems less noisy. So seems like the cleaning / electronics revamp did some good. 

Further characterization needs to be done, but the results of this test are encouraging. If we are able to get this kind of out of loop ALS noise with the IR beat, perhaps we can avoid having to frequently fine-tune the green beat alignment on the PSL table. It would also be ideal to mount this whole 1U setup in an electronics rack instead of leaving it on the PSL table.

Quote:

Photos + power budget + plan of action for using this box to characterize the green PDH locking to follow. 

GV Edit: I've added better photos to the 40m Google Photos page. I've also started a wiki page for this box / the proposed IR ALS  system. For the moment, all that is there is the datasheet to the Fiber Couplers used, I will populate this more as I further characterize the setup.

Attachment 1: IMG_7497.JPG  2.869 MB  Uploaded Fri Aug 25 17:00:54 2017  | Hide | Hide all
IMG_7497.JPG
Attachment 2: FOL_schematic.pdf  48 kB  Uploaded Fri Aug 25 17:11:22 2017  | Hide | Hide all
FOL_schematic.pdf
Attachment 3: 20170825_IR_ALS.pdf  73 kB  Uploaded Fri Aug 25 17:13:48 2017  | Hide | Hide all
20170825_IR_ALS.pdf
ELOG V3.1.3-